OpenStudy (anonymous):

in a certain year, the amount A of garbage in pounds produced afte t days by an average household is given by A=6.5t. How many days did it take for the average household to produce 150 pounds of garbarge?

OpenStudy (anonymous):

"How many days did it take for the average household to produce 150 pounds of garbarge?"\[ A = 150 \]We can substitute this into the equation already provided: \[ A=150 = 6.5t \]Now we just need to solve for \(t\)\[ 150 = 6.5t \]

OpenStudy (anonymous):

\[6.5t \]This is actually multiplication: \[ 6.5\times t \]The inverse is division \[ (6.5\times t)\div 6.5 = t \]If we do this to both sides of the equation simultaneously: \[ 150\div 6.5 = (6.5\times t)\div 6.5 \\ 150\div 6.5 = t \]

OpenStudy (anonymous):

23 days

OpenStudy (anonymous):

After 23 days, the average household will have produced just under 150 pounds of garbage. On the 24th day they have produced over 150 pounds.

OpenStudy (anonymous):

The real question is, do you think you could solve a similar problem on your own from now on?

OpenStudy (anonymous):

yeah now i just have to graph it

OpenStudy (anonymous):

how would you graph it if i am off a smudge it will be wrong